REINVESTIGATING THE Ã − X̃ SPECTRUM OF ETHYNYL RADICAL: NEW INSIGHTS INTO THE SPECTROSCOPY OF ^{2}Π −^{2} Σ^{+} VIBRONIC BANDS

\underline{ERIN N. SHARP-WILLIAMS}, MELANIE A. ROBERTS, DAVID J. NESBITT, JILA, National Institute of Standards and Technology and University of Colorado, and Department of Chemistry and Biochemistry, University of Colorado, Boulder, Colorado 80309; ROBERT F. CURL, Department of Chemistry and Rice Quantum Institute, Rice University, Houston, Texas 77005.

The ethynyl radical, C_{2}H, is an important reactive intermediate both in combustion processes, as it is readily formed in an acetylene (C_{2}H_{2}) flame, and in the chemistry of the interstellar medium, where it is suspected to be involved in the formation of longer carbon chain species (C_{n}H). We have recently interrogated several of the vibronic transitions to the lowlying excited Ã state from the vibrationless level of the ground electronic state via high resolution infrared spectroscopy. This was done using direct absorption laser spectroscopy in a slit-jet discharge supersonic expansion of C_{2}H_{2} diluted in a Neon/Helium gas mixture. In comparing our spectra with those already published using magnetic rotation spectroscopy^{a}^{,}^{b}, we find discrepancies between rovibronic frequencies in the 0_{0}^{0} band at 3600 cm^{−1}. The inconsistency is localized in the excited state, by comparison of 2-line combination differences with mm-wave measurements of the ground state by Thaddeus and coworkers^{c}. Calculating the ^{2}Π energy levels using both Hund’s case (a) and (b) basis sets and revisiting the analysis in the aforementioned work, we have determined that the discrepancies arise from a parity mislabeling of the lambda-doubled excited states. The improved low J signal intensities and resolution of satellite transitions that are observable under sub-Doppler, jet-cooled conditions complement the previous data and permit refinement of the rotational, spin-rotational, and lambda-doubling constants.

aW. B. Yan, C. B. Dane, D. Zeitz, J. L. Hall and R. F. Curl, J. Mol. Spectrosc. 123, 486 (1987).

bR. F. Curl, P. G. Carrick and A. J. Merer, J. Chem. Phys. 82, 3479 (1985).

cT. C. Killian, C. A. Gottlieb and P. Thaddeus, J. Chem. Phys. 127, 114320 (2007).